OpenStudy (anonymous):

A person who weighs 800N steps onto a scale that is on the floor of an elevated car. if the elevator accelerates upward at a rate of 5m/s^2, what will the scale read ?

OpenStudy (anonymous):

http://i.imgur.com/DhOCd14.jpg if you see image in the link, you'll see the illustration of the scenario. Newton's second law explains our scenario here.

OpenStudy (anonymous):

The R in this case is the Reaction/contact force exerted by the scale. Since the car is elevating upward, R will be greater than the weight of the person. So that's why: R > W where W = m*g m = W/g m = 800/9.8 m = 81.6 kg you have to understand that basically what Newton's second means is that when an object is experiencing various forces, and if there is a resultant force/Net force, then the object will accelerate towards a direction as well as magnitude of the Net force. since the whole car along with the person on the scale is moving upwards, it clearly means that the force exerted by the elevator is greater than the weight of the guy; this itself is the contact force which you see as a reading on the scale. So to set up the equation, it should look something like this: F = R-W ma = R - mg R = ma + mg R = (81.6*5) + 800 R = 1208.1 N so the reading on the scale will be 1208.1 N